# Colorful circuits and colorful topes in oriented matroids

### v1

@@ 1. Introduction
One of the most beautiful theorems in discrete geometry is the colorful Carathéodory theorem by Bárány [2]. It states: Given $d+1$ sets of points in $\mathbb{R}^{d}$, each of them containing the origin in its convex hull, there exists a way to pick one point from each set so that the resulting set still contains the origin in its convex hull. A generalization for oriented matroids has been proved by Holmsen [7] via a generalization of the topological colorful Helly theorem by Kalai and Meshulam [8]. Actually, the result by Holmsen (as well as that by Kalai and Meshulam) is a generalization of a much stronger version of the colorful Carathéodory theorem, where the colors are replaced by some (non-oriented) matroid constraint, and where the condition is less demanding. Yet, there exists a “conic” version of the colorful Carathéodory theorem: Given $d+1$ sets of vectors in $\mathbb{R}^{d+1}$, and a vector $p$ in $\mathbb{R}^{d+1}$ contained in the conic hull of each of these sets, there exists a way to pick one vector from each set so that the resulting set still contains $p$ in its conic hull. This conic version is also stronger than the original convex version.
@@
Let $C_{1},\dots,C_{r}$ be positive circuits of an oriented matroid of rank $r$ that all contain a common element $e$. Then there exist elements $f_{i}\in C_{i}\setminus\{e\}$  for $i=1,\dots,r$ such that the set $\{e,f_{1},\dots,f_{r}\}$ contains a positive circuit with $e$ as element.
@@
Let $C_{1},\dots,C_{r}$ be positive circuits of an oriented matroid of rank r1r-1. Then there exist elements $f_{i}\in C_{i}$  for $i=1,\dots,r$ such that the set $\{f_{1},\dots,f_{r}\}$ contains a positive circuit.
@@
Let $\sigma$ be a localization of the dual oriented matroid. We choose $\sigma$ in such a way that $\sigma(C)=+$ for every positive circuit of the original oriented matroid. This is possible thank to a result by Las Vergnas [4, Proposition 7.2.4] (applied with all elements of the ground set; this result requires to specify the signs taken by the localization on those elements via a vector $\alpha$,” which we set to the all-++ vector). Apply then Theorem 1.1 to the circuits $(C_{i},\sigma(C_{i}))$, with $e$ being the new element. This leads to elements $f_{i}\in C_{i}$ such that $\{e,f_{1},\ldots,f_{r}\}$ contains a positive circuit of the lifted oriented matroid. Contracting $e$ shows that $\{f_{1},\ldots,f_{r}\}$ contains a positive circuit of the original oriented matroid. ∎
@@
Consider a uniform oriented matroid on $[n]$. For every collection of $n$ topes $T_{1},T_{2},\ldots,T_{n}$ that agree on element $n$, there exists a tope $\widetilde{T}$ agreeing with each $T_{i}$ on a distinct element.
@@
Given $n$ topes $T_{1},\dots,T_{n}$ of a uniform oriented matroid on $n$ elements, under what conditions does there exist a tope $\widetilde{T}$ agreeing with each tope $T_{i}$ on a distinct element?
@@
Consider a uniform oriented matroid of rank $r$ on $[n]$. For every collection of topes $T_{1},T_{2},\ldots,T_{r}$ and positive integers $n_{1},n_{2},\ldots,n_{r}$ summing up to $n$, there exist a tope $\widetilde{T}$ and a partition $n]=J_{1}\cup J_{2}\cup\cdots\cup J_{r}$ such that $|J_{i}|=n_{i}$ and $\widetilde{T}(j)=T_{i}(j)$ for all $i\in[r]$ and $j\in J_{i}$.
@@
Consider a uniform oriented matroid of rank two on $n$ elements. For every collection of $n$ topes $T_{1},T_{2},\ldots,T_{n}$, there exists a tope $\widetilde{T}$ agreeing with each $T_{i}$ on a distinct element.
@@ 2.2. Preliminary results
In this subsection, we consider an oriented matroid $\mathcal{M}$ on $[n]$. We establish and present elementary results on $\mathcal{V}$ and $\mathcal{L}$—its sets of vectors and covectors—seen as partially ordered sets (with the usual order of oriented matroids). The proofs will rely on the topological representation theorem; see [4, Chapter 5]. We emphasize that because of the choice we made regarding the statements of the theorems—with circuits for Theorem 1.1 and with topes for Theorems 1.3 and 1.4—the topological representation theorem will be used sometimes for $\mathcal{M}$ and sometimes for its dual.
@@
We take the barycentric subdivision of $\partial\sigma$, whose existence is guaranteed by $X$ being a regular cell complex. We denote by $\operatorname{sd}(\partial\sigma)$ this barycentric subdivision. Consider the collection of stars $\operatorname{st}(u)$ for $u$ ranging $U^{0}$. Here, we consider “closed stars”, i.e., the collection of simplices $\tau$ in $\operatorname{sd}(\partial\sigma$) that together with $u$ form a simplex of $\operatorname{sd}(\partial\sigma)$. We make now two crucial remarks. The first one is that these stars are simplicial subcomplexes. The second one goes as follows: for every subset $U^{\prime}\subseteq U^{0}$, either there is no cell of $\partial\sigma$ containing $U^{\prime}$ completely, in which case $\bigcap_{u\in U^{\prime}}\operatorname{st}(u)=\varnothing$, or there is a (unique) minimal cell $\tau$ containing $U^{\prime}$, in which case $\bigcap_{u\in U^{\prime}}\operatorname{st}(u)$ has all its maximal simplices containing the barycenter of $\tau$. (Note that we use here the lattice structure of the face poset of $\sigma$.) In other words, if $\bigcap_{u\in U^{\prime}}\operatorname{st}(u)$ is non-empty, it is a cone with apex the barycenter of $\tau$, and thus contractible. These two remarks make that we can apply the nerve theorem, in a version for simplicial subcomplexes [3, Theorem 6]: homology of $\bigcup_{u\in U^{0}}\operatorname{st}(u)$ is that of the nerve of the stars $\operatorname{st}(u)$ for $u$ ranging over $U^{0}$. Since $\bigcap_{u\in U^{0}}\operatorname{st}(u)$ contains the barycenter of the cell $\sigma\cap S_{e}$, it is non-empty, which means that the nerve of the stars is a full simplex, and this implies that $\bigcup_{u\in U^{0}}\operatorname{st}(u)$ is acyclic. We finish by the following chain of equalities and homotopy equivalence. We denote by $B$ the vertex set of $\operatorname{sd}(\partial\sigma)$ and by $B^{0}$ its subset corresponding to barycenters of cells hitting $U^{0}$:
@@
It is homologically (|I|2)(|I|-2)-connected by Lemma 2.3, applied to the oriented matroid restricted to the ground set $\bigcup_{i\in I}\underline{C}_{i}$, with $h=|I|$. Define a labeling $\lambda$ of the positive vectors of $\mathcal{M}$ by elements from $[r]$ as follows: for a positive vector $X$, let $\lambda(X)$ be the index $i\in[r]$ such that $X$ selects the maximum number of elements of $C_{i}$; in case of a tie, choose the smallest possible $i$. This definition makes that $\lambda(X)\in I$ when $X\in\mathsf{K}^{I}$. (Here, we use that $e$ is not a loop.) The family $\mathsf{K}^{I}$ and the labeling $\lambda$ satisfy the conditions of Lemma 2.1. Hence, there exist positive vectors $X_{1}\prec X_{2}\prec\cdots\prec X_{r}$ such that $\lambda(X_{i})$ get distinct values. By construction, the height of $X_{r}$ is at most $r$, which implies by Proposition 2.2 that $|\underline{X}_{r}|\leqslant 2r$ (Note that here we use $\operatorname{rk}(\mathcal{M})\leqslant r$.) The element $e$ is in the support of all $X_{i}$ by definition of the $\mathsf{K}^{I}$. Thus, there is an $i^{\star}$ such that circuit $C_{i^{\star}}$ contributes in $X_{r}$ by at most one element (apart from $e$). There is a vector $X_{j}$ whose image by $\lambda$ is $i^{\star}$. Necessarily, no circuit $C_{i}$ contributes by more than one to this vector $X_{j}$, and since $X_{1}\preccurlyeq X_{j}$, no circuit $C_{i}$ contributes by more than one to $X_{1}$. This latter vector $X_{1}$ being of height $1$, it is a positive circuit that shares at most one element (outside $e$) with each $C_{i}$, as desired. ∎
@@
Consider a uniform oriented matroid $\mathcal{M}$ on $[n]$ as in the statement of these theorems. Let $T_{1},T_{2},\ldots,T_{k}$ be topes. For a non-empty subset $I$ of $[k]$, we set
@@
Since every vertex of $\tau$ corresponds to a tope, repeated application of vector elimination, in its strong version by Edmonds, Fukuda, and Mandel [6, 9] (we use actually [4, Corollary 3.7.9, Chapter 3]), shows that this $X$ is a covector of $\mathcal{M}$. Set $f(\tau)\coloneqq X$. This defines an order-reversing map $f\colon\Delta(\mathsf{L}^{I})\to\mathcal{L}_{J^{+},J^{-}}$. Given $Y\in\mathcal{L}_{J^{+},J^{-}}$, define $\tau^{Y}=\tau_{1}^{Y}\times\tau_{2}^{Y}\times\cdots\times\tau_{n}^{Y}$ by setting $\tau_{j}^{Y}\coloneqq\sigma^{I}$ if $Y(j)=0$ and by $\tau_{j}^{Y}\coloneqq\sigma^{\{i\in I\colon T_{i}(j)=Y(j)\}}$ otherwise. The oriented matroid $\mathcal{M}$ being uniform, $\tau^{Y}$ is a face of $\mathsf{L}^{I}$. The faces $\tau$ of $\mathsf{L}^{I}$ such that $f(\tau)\succcurlyeq Y$ are precisely the faces of $\tau^{Y}$. Since this latter is a polytope, it is contractible. The desired result follows then from the Quillen lemma (Lemma 2.5). ∎
@@
In this proof, we work with $k=r=\operatorname{rk}(\mathcal{M})$. Let $\tau=\tau_{1}\times\cdots\times\tau_{n}$ be a face of $\mathsf{L}^{[r]}$ of dimension at most r1r-1. Define $\lambda(\tau)$ as the smallest index $i$ such that there are at least $n_{i}$ indices $j$ with $\tau_{j}=\{i\}$. Such an $i$ exists because the dimension of $\tau$ implies that there are at least $n-r+1>\sum_{i=1}^{r}(n_{i}-1)$ indices $j$ such that $\tau_{j}$ is a singleton. Thanks to Lemma 2.6 and by the definition of $\lambda$, the condition of Lemma 2.1 is satisfied for $\mathsf{K}^{I}$ being the barycentric subdivision of the (|I|1)(|I|-1)-skeleton of $\mathsf{L}^{I}$. This means that there exist simplices $\tau^{1}\subset\cdots\subset\tau^{r}$ of $\mathsf{L}^{[r]}$ such that $\lambda(\tau^{i})=i$ for every $i\in[r]$, with $\tau^{1}$ of dimension 0. Setting $J_{i}$ as the indices $j$ such that $\tau_{j}^{1}=\{i\}$ (where $\tau^{1}=\tau_{1}^{1}\times\cdots\times\tau_{r}^{1}$), we get the desired partition. ∎

### v2

@@ 1. Introduction
One of the most beautiful theorems in discrete geometry is the colorful Carathéodory theorem by Bárány [2]. It states: Given $d+1$ sets of points in $\mathbb{R}^{d}$, each of them containing the origin in its convex hull, there exists a way to select one point from each set so that the resulting set still contains the origin in its convex hull. A generalization for oriented matroids has been proved by Holmsen [7] via a generalization of the topological colorful Helly theorem by Kalai and Meshulam [8]. Actually, the result by Holmsen (as well as that by Kalai and Meshulam) is a generalization of a much stronger version of the colorful Carathéodory theorem, where the colors are replaced by some (non-oriented) matroid constraint, and where the condition is less demanding. Yet, there exists a “conic” version of the colorful Carathéodory theorem: Given $d+1$ sets of vectors in $\mathbb{R}^{d+1}$, and a vector $p$ in $\mathbb{R}^{d+1}$ contained in the conic hull of each of these sets, there exists a way to select one vector from each set so that the resulting set still contains $p$ in its conic hull. This conic version is also stronger than the original convex version.
@@
Consider an oriented matroid of rank $r$. For every collection of $r$ positive circuits $C_{1},\dots,C_{r}$ that share a common element $e$, there is a way to select at most one element from each $C_{i}\setminus\{e\}$ so as to form with $e$ a positive circuit.
We emphasize that a $C_{i}$ may share several elements with the resulting positive circuit. A very formal statement could finish by “. . . there exist a positive circuit $\widetilde{C}$ containing $e$ and an injective map $\pi\colon\widetilde{C}\setminus\{e\}\to[r]$ such that $f\in C_{\pi(f)}$ for all $f\in\widetilde{C}\setminus\{e\}$.
@@
Consider an oriented matroid of rank r1r-1. For every collection of $r$ positive circuits $C_{1},\dots,C_{r}$, there is a way to select at most one element from each $C_{i}$ so as to form a positive circuit.
@@
Let $\sigma$ be a localization of the dual oriented matroid. We choose $\sigma$ in such a way that $\sigma(C)=+$ for every positive circuit of the original oriented matroid. This is possible thank to a result by Las Vergnas [4, Proposition 7.2.4] (applied with all elements of the ground set; this result requires to specify the signs taken by the localization on those elements via a vector $\alpha$,” which we set to the all-++ vector). Apply then Theorem 1.1 to the circuits $(C_{i},\sigma(C_{i}))$, with $e$ being the new element. This shows that there is a way to select at most one element from each $C_{i}$ so as to form with $e$ a positive circuit of the lifted oriented matroid. Contracting $e$ provides the desired positive circuit of the original oriented matroid. ∎
@@
Consider a uniform oriented matroid with $n$ elements. For every collection of $n$ topes $T_{1},\dots,T_{n}$ that agree on some element, there exists a tope $\widetilde{T}$ agreeing with each $T_{i}$ on a distinct element.
We emphasize that, in general, the tope $\widetilde{T}$ does not necessarily belong to the collection of the $n$ topes.
@@
Given a collection of $n$ topes $T_{1},\dots,T_{n}$ of a uniform oriented matroid on $n$ elements, under what conditions does there exist a tope $\widetilde{T}$ agreeing with each tope $T_{i}$ on a distinct element?
@@
Consider a uniform oriented matroid of rank $r$ and with $n$ elements. For every collection of $n$ topes $T_{1},\dots,T_{n}$ among which at most $r$ are pairwise distinct, there exists a tope $\widetilde{T}$ agreeing with each $T_{i}$ on a distinct element.
An alternative and more formal way of stating this theorem could be: Consider a uniform oriented matroid of rank $r$ on $[n]$. For every collection of topes $T_{1},\ldots,T_{r}$ and positive integers $n_{1},\ldots,n_{r}$ summing up to $n$, there exist a tope $\widetilde{T}$ and a partition $n]=J_{1}\cup\cdots\cup J_{r}$ such that $|J_{i}|=n_{i}$ and $\widetilde{T}(j)=T_{i}(j)$ for all $i\in[r]$ and $j\in J_{i}$. The original $n$ topes are obtained by considering $n_{i}$ copies of each $T_{i}$.
@@
Consider a uniform oriented matroid of rank two on $n$ elements. For every collection of $n$ topes $T_{1},\ldots,T_{n}$, there exists a tope $\widetilde{T}$ agreeing with each $T_{i}$ on a distinct element.
@@ 2.1. A “Sperner–Meshulam” lemma
In this proof, we use chains and chain maps as defined by Munkres [12, Chapter 1, Sections 5 and 12]. In particular, degenerate simplices are not considered.
@@ 2.2. Preliminary results
In this subsection, we consider an oriented matroid $\mathcal{M}$ on $[n]$. We establish and present elementary results on $\mathcal{V}$ and $\mathcal{L}$—its sets of vectors and covectors—seen as partially ordered sets (with the usual order on vectors and covectors of an oriented matroid). The proofs will rely on the topological representation theorem; see [4, Chapter 5]. We emphasize that because of the choice we made regarding the statements of the theorems—with circuits for Theorem 1.1 and with topes for Theorems 1.3 and 1.4—the topological representation theorem will be used sometimes for $\mathcal{M}$ and sometimes for its dual.
@@
We take the barycentric subdivision of $\partial\sigma$, whose existence is guaranteed by $X$ being a regular cell complex. We denote by $\operatorname{sd}(\partial\sigma)$ this barycentric subdivision. Consider the collection of stars $\operatorname{st}(u)$ for $u$ ranging $U^{0}$. Here, we consider “closed stars,” i.e., the collection of simplices $\tau$ in $\operatorname{sd}(\partial\sigma$) that together with $u$ form a simplex of $\operatorname{sd}(\partial\sigma)$. We make now two crucial remarks. The first one is that these stars are simplicial subcomplexes. The second one goes as follows: for every non-empty subset $U^{\prime}\subseteq U^{0}$, there is a (unique) minimal cell $\tau$ containing $U^{\prime}$ and $\bigcap_{u\in U^{\prime}}\operatorname{st}(u)$ has all its maximal simplices containing the barycenter of $\tau$. (Note that we use here the lattice structure of the face poset of $\sigma$.) In other words, $\bigcap_{u\in U^{\prime}}\operatorname{st}(u)$ is always non-empty, it is a cone with apex the barycenter of $\tau$, and it is thus contractible. These two remarks make that we can apply the nerve theorem, in a version for simplicial subcomplexes [3, Theorem 6]: homology of $\bigcup_{u\in U^{0}}\operatorname{st}(u)$ is that of the nerve of the stars $\operatorname{st}(u)$ for $u$ ranging over $U^{0}$. Since $\bigcap_{u\in U^{0}}\operatorname{st}(u)$ contains the barycenter of the cell $\sigma\cap S_{e}$, it is non-empty, which means that the nerve of the stars is a full simplex, and this implies that $\bigcup_{u\in U^{0}}\operatorname{st}(u)$ is acyclic. We finish by the following chain of equalities and homotopy equivalence. We denote by $B$ the vertex set of $\operatorname{sd}(\partial\sigma)$ and by $B^{0}$ its subset corresponding to barycenters of cells hitting $U^{0}$:
@@
It is homologically (|I|2)(|I|-2)-connected by Lemma 2.3, applied to the oriented matroid restricted to the ground set $\bigcup_{i\in I}\underline{C}_{i}$, with $h=|I|$. Define a labeling $\lambda$ of the non-zero positive vectors of $\mathcal{M}$ by elements from $[r]$ as follows: for a non-zero positive vector $X$, let $\lambda(X)$ be the index $i\in[r]$ such that $X$ selects the maximum number of elements of $C_{i}$; in case of a tie, choose the smallest possible $i$. This definition makes that $\lambda(X)\in I$ when $X\in\mathsf{K}^{I}$. (Here, we use that $e$ is not a loop.) The family $\mathsf{K}^{I}$ and the labeling $\lambda$ satisfy the conditions of Lemma 2.1. Hence, there exist positive vectors $X_{1}\prec\cdots\prec X_{r}$ such that $\lambda(X_{i})$ get distinct values. By construction, the height of $X_{r}$ is at most $r$, which implies by Proposition 2.2 that $|\underline{X}_{r}|\leqslant 2r$ (Note that here we use $\operatorname{rk}(\mathcal{M})\leqslant r$.) The element $e$ is in the support of all $X_{i}$ by definition of the $\mathsf{K}^{I}$. Thus, there is an $i^{\star}$ such that circuit $C_{i^{\star}}$ contributes in $X_{r}$ by at most one element (apart from $e$). There is a vector $X_{j}$ whose image by $\lambda$ is $i^{\star}$. Necessarily, no circuit $C_{i}$ contributes by more than one to this vector $X_{j}$, and since $X_{1}\preccurlyeq X_{j}$, no circuit $C_{i}$ contributes by more than one to $X_{1}$. This latter vector $X_{1}$ being of height $1$, it is a positive circuit whose elements outside $e$ can be obtained by selecting at most one element from each $C_{i}\setminus\{e\}$.∎
@@
Consider a uniform oriented matroid $\mathcal{M}$ on $[n]$ as in the statement of these theorems. Let $T_{1},\ldots,T_{k}$ be topes. For a non-empty subset $I$ of $[k]$, we set
@@
Since every vertex of $\tau$ corresponds to a tope, repeated application of vector elimination, in its strong version by Edmonds, Fukuda, and Mandel [6, 9] (we use actually [4, Corollary 3.7.9, Chapter 3]), shows that this $X$ is a covector of $\mathcal{M}$. Set $f(\tau)\coloneqq X$. This defines an order-reversing map $f\colon\Delta(\mathsf{L}^{I})\to\mathcal{L}_{J^{+},J^{-}}$, where $\Delta(\mathsf{L}^{I})$ is the order complex of $\mathsf{L}^{I}$. Given $Y\in\mathcal{L}_{J^{+},J^{-}}$, define $\tau^{Y}=\tau_{1}^{Y}\times\tau_{2}^{Y}\times\cdots\times\tau_{n}^{Y}$ by setting $\tau_{j}^{Y}\coloneqq\sigma^{I}$ if $Y(j)=0$ and by $\tau_{j}^{Y}\coloneqq\sigma^{\{i\in I\colon T_{i}(j)=Y(j)\}}$ otherwise. The oriented matroid $\mathcal{M}$ being uniform, $\tau^{Y}$ is a face of $\mathsf{L}^{I}$. The faces $\tau$ of $\mathsf{L}^{I}$ such that $f(\tau)\succcurlyeq Y$ are precisely the faces of $\tau^{Y}$. Since this latter is a polytope, it is contractible. The desired result follows then from the Quillen lemma (Lemma 2.5). ∎
@@
In this proof, we work with $k=r=\operatorname{rk}(\mathcal{M})$, and we actually prove the more formal statement given just after Theorem 1.4: we consider $r$ topes $T_{1},\ldots,T_{r}$, not necessarily distinct, and positive integers $n_{1},\ldots,n_{r}$ summing up to $n$. Let $\tau=\tau_{1}\times\cdots\times\tau_{n}$ be a face of $\mathsf{L}^{[r]}$ of dimension at most r1r-1. Define $\lambda(\tau)$ as the smallest index $i$ such that there are at least $n_{i}$ indices $j$ with $\tau_{j}=\{i\}$. Such an $i$ exists because the dimension of $\tau$ implies that there are at least $n-r+1>\sum_{i=1}^{r}(n_{i}-1)$ indices $j$ such that $\tau_{j}$ is a singleton. Thanks to Lemma 2.6 and by the definition of $\lambda$, the condition of Lemma 2.1 is satisfied for $\mathsf{K}^{I}$ being the barycentric subdivision of the (|I|1)(|I|-1)-skeleton of $\mathsf{L}^{I}$. This means that there exist simplices $\tau^{1}\subset\cdots\subset\tau^{r}$ of $\mathsf{L}^{[r]}$ such that $\lambda(\tau^{i})=i$ for every $i\in[r]$, with $\tau^{1}$ of dimension 0. Setting $J_{i}$ as the indices $j$ such that $\tau_{j}^{1}=\{i\}$ (where $\tau^{1}=\tau_{1}^{1}\times\cdots\times\tau_{r}^{1}$), we get the desired partition. ∎